OpenStudy (anonymous):

Can anyone help! 1. Simplify each equation. show all work, step-by-step. a. sqrt6w^3 * sqrt12w b. (125)^-2\3 2.Simplify each expression. Rationalize denominators. Show all work, step-by-step. a.10/sqrt5 b.sqrt w^4/x^3

OpenStudy (abb0t):

\[\frac{ a^b }{ a^g } = a^{b-g}\] Similarly: \[\frac{ 1 }{ \sqrt{a} } = a^{- \frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

How would I use that?

OpenStudy (abb0t):

Well, you would use that for the 1st and last question.

OpenStudy (abb0t):

It might also help if you know that \[\frac{ 1 }{ a^2 } = a^{-2}\]

OpenStudy (anonymous):

so would the sqrt 6w^3 go over sqrt 12w

OpenStudy (anonymous):

1.a)\[\sqrt{6w^3}\times \sqrt{12w}\] \[=w \sqrt{6w}\times2\sqrt{3w}\] \[=2w \sqrt{18w^2}\] \[=6w^2\sqrt{2}\] b)\[(125)^{-\frac{ 2 }{ 3 }}\] \[=\frac{ 1 }{ \sqrt[3]{(125)^2} }\] \[=\frac{ 1 }{ 25 }\]

OpenStudy (anonymous):

2. a)\[\frac{ 10 }{\sqrt5 }\times \frac{ \sqrt5 }{ \sqrt5 }\] \[=\frac{ 10\sqrt5 }{ 5 }\] \[=2\sqrt5\]

OpenStudy (anonymous):

b)\[\sqrt{\frac{ w^4 }{ x^3 }}\] \[=\frac{ w }{ x }\sqrt{\frac{ w^2 }{ x }}\] \[=\frac{ w \sqrt{w^2} }{ x \sqrt{x} }\times \frac{ \sqrt{x} }{ \sqrt{x} }\] \[=\frac{ w^2 \sqrt{x} }{ x^2 }\]